If a simple harmonic oscillator has velocity function of:

v                =      100πAsin(100πt – 0.75π)   Find:

         i) an expression for displacement, x.    

         ii) an expression for acceleration, a. 


1
Expert's answer
2021-11-29T11:45:10-0500

Given:

"v =100πA\sin(100πt – 0.75π)"


(i) the displacement is related with velocity by formula

"x=\int vdt\\=100πA\int\sin(100πt – 0.75π)\\=-A\cos(100πt – 0.75π)"

(ii) the acceleration

"a=\frac{dv}{dt}\\
=(100πA\sin(100πt – 0.75π) )'\\
=100^2\pi^2A\cos(100πt – 0.75π)"